OpenStudy (anonymous):

HELP!!! 2/3t = 20 WHAT IS T????

OpenStudy (anonymous):

\[2/3 * t = 20/1 \] so that means that if you rearrange it you get \[2*t = 3*20 = 60\] and then divide both sides by 2 which gives: \[t = 60/2 = 30\]